### v1

@@ Abstract
We report the performance of a magnetically silent optically pumped cesium magnetometer with a statistical sensitivity of 3.9 pT and a stability of 90 fT over 150 seconds of measurement. Optical pumping with coherent, linearly-polarized, resonant light leads to a relatively long-lived polarized ground state of the cesium vapour contained in a measurement cell. The state precesses at its Larmor frequency in the magnetic field to be measured. Nonlinear magneto-optical rotation then leads to the rotation of the plane of polarization of a linearly polarized probe laser beam. The rotation angle is modulated at twice the Larmor frequency. A measurement of this frequency constitutes an absolute measurement of the magnetic field magnitude. Featuring purely optical operation, non-magnetic construction, low noise floor, and high stability, this sensor will be used for the upcoming TUCAN electric dipole moment experiment and other highly sensitive magnetic applications.
@@
magnetometer, magnetometry, neutron electric dipole moment, all-optical, alkali, atomic vapour, magnetically silent, nonlinear magneto-optical rotation, Faraday rotation, Bell-Bloom, Cesium, free spin precession
@@ 1 Introduction
Measurements of the nEDM require extremely well measured and controlled magnetic fields [20, 21, 14]. Cs magnetometry was used in conjunction with Hg comagnetometry in the latest upper limit paper [14]. The Cs sensors used by the PSI nEDM collaboration were demonstrated to have sensitivities of 0.75-8 pT/$\sqrt{\mathrm{Hz}}$ and absolute accuracies of 45-90 pT per measurement [20]. However, due to the method by which these sensors are operated they are not magnetically silent (see section 2), and can therefore not be used while neutrons are being measured and thus have limited utility.
@@
In this work we demonstrate a statistical sensitivity of 3.9 pT and a stability of 90 fT over 150 seconds of measurement in a configuration that operates purely optically, using NMOR to both create and probe a polarized atomic state. This allows precision magnetic field measurements to be done during the EDM experiment without perturbing the neutrons [25].
@@ 2 Configuration
Figure 1: The beam geometry for this style of magnetometry. Both the pump and probe beams are linearly polarized, and the electric field vectors of the two light fields are aligned. The probe beam propagates through the vapor cell along the magnetic field vector to be measured, while the pump can be anywhere in the plane formed by the two beams in this figure. The magnetic field information is encoded in the time dependence of the polarization angle $\phi$ of the probe beam after it exits the vapour cell
@@
The Bell-Bloom style FSP configuration used in our system also measures the Larmor frequency of Cs to deduce the magnetic field experienced by the atoms; however, it uses nonlinear magneto-optical rotation [26, 29] rather than absorption to both polarize the atoms and interrogate their Larmor frequency. In this configuration, the electric field vector of both the pump and probe beams are perpendicular to the magnetic field, as in figure 1. When the pump beam is turned on, atoms in the beam are driven into a coherent non-interacting state, polarizing the atoms along the electric field vector of the pump light. Since we are using linearly polarized light, the polarization is an aligned state rather than the oriented state used by the $M_{x}$ configuration. Once they are driven into non-interacting states, the atoms immediately start to precess in the magnetic field, rotating them towards being interacting states. So as to not de-pump these atoms, the pump beam is then quickly turned off again. After half of one Larmor period the previously polarized atoms are again non-interacting, and the pump beam is switched back on, creating another set of polarized atoms in phase with the previously created atoms. By repeating this process, a large population of such polarized atoms can be produced. Once the population is maximized, the pump beam is switched off. A separate, much weaker probe beam is then passed through the atoms, propagating parallel to the magnetic field. The polarized state of the Cs exhibits an axis of birefringence that precesses along with the atoms. This causes the angle of the linearly polarized probe light, $\phi$, to be modulated at twice the Larmor frequency of the atoms because of the twofold symmetry of the aligned state. This modulation is detected via polarimetry after the probe beam exits the vapour.
@@
The FSP configuration eliminates the issue of RF crosstalk and interference by eliminating the use of RF entirely. The issue of phase drift is eliminated by running exclusively in FSP mode, while the issue of the vector light shift is reduced by using linearly polarized light for the probe beam. Additionally, servicing all sensors with a single probe laser means any systematic shifts associated with laser diode intensity or frequency fluctuations should not impact relative measurements between sensors, meaning gradient extraction is not affected. These same changes are planned by the PSI collaboration for their upcoming n2EDM experiment [2] and have been implemented by the panEDM collaboration as well [1, 30, 31, 32].
@@ 3 Apparatus
All fibre in this system before the sensor is either polarization maintaining (PM) or polarizing (PZ) fibre. This ensures very clean linear polarization in the sensor head. The pump laser is launched into the input fibre of a fibre-coupled acousto-optic modulator (AOM). This is then attached to a $1\times 8$ fibre splitter (also using PM fibre) which splits the pump light equally into 8 pigtailed PM fibres, each of which can provide pump light to a sensor. The probe light is launched directly into another $1\times 8$ PM splitter. The overall intensity of the probe beams are controlled by a linear polarizer and half-wave plate in the free space section of the beam. Individual ports can be additionally attenuated using an adjustment screw on the body of the splitter. The free space beams are $\sim$100 mW, typical launch efficiencies with the current opto-mechanical setup are 20-30%, leading to $\sim$1-2 mW available for use in the sensor heads after accounting for all losses from splitting and attenuation. This is sufficient power to optimize the pump light [25] and more than enough to provide probe power. The pump beams can be globally attenuated by adjusting the high level of the square wave operating the AOM. The probe beams are typically $<5~{}\mu$W in the sensor head after being adjusted using screw-attenuators built into the fibre splitter output ports. $<5~{}\mu$W is enough to reach shot-noise-limited performance in the photodiodes without limiting the coherence time of the Cs via de-pumping. The exact power is optimized to minimize the uncertainty in the frequency fit of the FSP signal.
The magnetically sensitive component of these sensors is an evacuated glass cell with a small amount of Cs held in a reservoir, attached to the cell via a capillary tube. This cell was manufactured by Precision Glass Blowing (PGB) [33]. The cell is treated with an anti-relaxation alkene coating available from PGB prior to being filled, which greatly extends the polarization lifetime of the atoms in the cell by reducing polarization-destroying wall collisions. Light is delivered to the Cs cell where it is collimated into a free-space beam to pass through the cell and interact with the Cs. After passing though the Cs cell the probe beam is directed through a polarizing beamsplitter and the resulting two beams are launched into MM fibre to be carried out of the magnetically sensitive environment for detection and analysis. This sensor head was designed and manufactured by Southwest Sciences [34] in consultation with W. Klassen and J. Martin, and incorporates design improvements developed in Munich [1, 30, 31, 32].
@@
The optical components are held in a plastic clamshell with a very similar design to the current panEDM Cs design [30], 3D printed in undyed ABS plastic. Optical components are cemented to one another at the optical interfaces using RD3-74 epoxy [35]. The epoxy provides the structural connection and maintains alignment between optical components, with the clamshell providing a mounting point for the whole sensor. Linearly polarized light from two laser diodes is delivered to the sensor head via polarizing fibre optics. A gradient-index (GRIN) lens launches light from the probe fibre to a $\sim$1 mm2 free-space beam, which is then directed through a clean-up linear polarizer, followed by the vapour cell via internal reflection through a right-angled prism. This probe launching assembly is mounted to the main ABS body using a coupling which constrains the motion of the probe assembly to linear motion in a plane, which helps with alignment. Linear translation stages are then used to position the assembly to maximize coupling into the exit fibers prior to cementing the prism to the face of the vapour cell as the final step in optical alignment. The beam is then again redirected via a right-angled prism through a Rochon walk-off polarizing beam splitter which splits the beam into orthogonal, linearly polarized components. These components are then launched back into fibre optics: much larger core MM fibres are used on this side so as to collect the most light possible. The two polarization components are then used for polarimetry of the probe light. A GRIN lens also launches linearly polarized pump light into the cell after passing it through a clean-up polarizer, due to the geometry of our setup it can pass straight through the cell at 90 to the probe beam, since the electric field vectors of the two beams are aligned. This choice of beam geometry was largely driven by considering ease of manufacture, launching both beams into the end of the vapour cell would be difficult to do. Any loss in pumping efficiency due to the shorter optical path length can be compensated with pump power.
@@
The Rochon prism is oriented with its fast-axis at a 45 angle relative to the probe beam polarization axis so that it evenly splits the unperturbed probe light into two components of equal power. This means that for small rotations, the degree of rotation of polarized light is directly proportional to the difference in power between the two components. The MM fibres are plugged into a polarimetry board, which amplifies the photo-current difference from the two photodiodes watching the two fibres. The resulting signal is then captured on an oscilloscope, or via a 32 channel DAQ system made by D-tAqc [36]. The D-tAqc system has a bandwidth of 200 kHz with 16 bit resolution. The polarimetry board differencing circuit has a high-frequency cutoff of around 15 kHz, and is AC coupled since we are not interested in static imbalances in the two polarization channels. This relatively low frequency cut-off was due to regulatory requirements imposed on the manufacturer rather than any physics motivation.
@@
Figure 5: The amplified, differential photo-current corresponding to the degree of rotation of polarized light as a function of time during the polarization and the free spin precession phases of the Cs gas. Pumping begins at T=0 on this plot, and stops just after 40 ms when the signal reaches its maximum. During this time the pump beam is modulated into a 20% duty cycle square wave at $\sim$7 kHz, approximately the Larmor frequency of Cs in a 1 $\mu$T field. The remainder of the signal is the free spin precession portion which is later fit to an exponentially decaying sine wave. A single instance of this pump-probe cycle is sometimes referred to as one free spin precession cycle, or one “FSP” cycle
Figure 6: The measured magnetic field (blue line) and measured coil voltage (orange line) versus time. The voltage across is being monitored by an 8.5 digit multimeter from Keithley, which is set to an integration time of 4/60 of a second, or 4 power line cycles, matching the integration time of the FSP signal measuring the magnetic field
@@
To operate in 1 $\mu$T we use passive magnetic shielding to reduce both earth’s field and fluctuations in it. The shield is four nested cylinders of $\mu$-metal, a highly magnetically permeable alloy of steel. The outer cylinder is 40 cm long with a 15 cm radius. The innermost volume contains a 3D printed saddle-wound coil designed to provide a homogeneous transverse field inside the shield. The volume inside this coil is 20 cm long with radius 3.5 cm. The axial shielding factor has been measured [37] to be 1.4×$\times 10^{7}$ with all endcaps installed, and while the transverse shielding factor has not been measured it is expected to be much higher. Since this inner volume is quite small, the uniformity is on the order of nT/cm. This is quite non-uniform, but provides very high AC and DC shielding factors. All sensor tests are done in this shield.
@@ 4 Results
Figure 6 shows a magnetic field measurement with the system. After being split into individual FSPs the signal is filtered with a windowed-sinc FIR filter, with a pass-band from 6-8 kHz, transitional bands of 2 kHz on each side, and $<0.04$ dB of ripple in the pass band. The filtered signal is fit to an exponentially decaying sine wave
@@
where $A$ is the amplitude, $T_{2}$ is the time constant of the decay envelope, $\omega$ is the frequency of the precession, and $\Phi$ is the phase. $\omega$ is directly scaled to be the magnetic field strength via the gyromagnetic ratio of Cs. While this data was taken, the coil voltage was being monitored across a precision resistor using a digital multimeter (DMM) with 8.5 digits of precision. The voltage is scaled to a magnetic field via the ratio of the averages of the two data sets so they can be directly compared without a perfectly accurate coil constant. The Allan deviation of the same data is shown in Figure 8. The voltage monitor from the power supply is significantly less stable than the sensor readings for most integration times. We attribute this to environmental effects on both the DMM itself and the connection to the monitor port. The sensor’s Allan deviation reaches a minimum of 90$\pm$8 fT at an integration time of 150 seconds, so we can conclude that our sensors are at least this stable. The minimum integration time Allan deviation corresponds to the average deviation between successive points and can be taken as a measurement of sensitivity. In this case the sensitivity is 3.9 $\pm$ 0.1 pT.
Figure 9: Magnetic field measurements of two sensors (a and b) together with the difference between them (c) as a function of time. The sensors are operated simultaneously, and when the field is stepped up and down, the change in field is observed in each sensor individually. The sensor measurements are subtracted, as shown in (c), and the difference remains at their combined noise floor. Because of the relatively inhomogeneous field in the magnetic shield, the two sensors are operating at different fields and thus are not able to be perfectly optimized simultaneously given they are driven by the same optical system with a common pump cycle frequency. The higher noise floor is the result of this. The spike in the difference signal around 60 seconds can be associated with the changing magnetic field, and is most likely the result of the field changing partway through an FSP ringdown

### v2

@@ Abstract
We report the performance of a magnetically silent optically pumped cesium magnetometer with a statistical sensitivity of 3.5 pT/$\sqrt{\mathrm{Hz}}$ at 1 Hz and a stability of 90 fT over 150 seconds of measurement. Optical pumping with coherent, linearly-polarized, resonant light leads to a relatively long-lived polarized ground state of the cesium vapour contained in a measurement cell. The state precesses at its Larmor frequency in the magnetic field to be measured. Nonlinear magneto-optical rotation then leads to the rotation of the plane of polarization of a linearly polarized probe laser beam. The rotation angle is modulated at twice the Larmor frequency. A measurement of this frequency constitutes an absolute measurement of the magnetic field magnitude. Featuring purely optical operation, non-magnetic construction, low noise floor, and high stability, this sensor will be used for the upcoming TUCAN electric dipole moment experiment and other highly sensitive magnetic applications. Novel aspects of the system include commercial construction and the ability to operate up to 24 sensors on a single probe laser diode.
@@
magnetometer, magnetometry, neutron electric dipole moment, all-optical, alkali, atomic vapour, magnetically silent, nonlinear magneto-optical rotation, Faraday rotation, Bell-Bloom, Cesium, free precession decay
@@ 1 Introduction
Measurements of the nEDM require extremely well measured and controlled magnetic fields [20, 21, 14]. Cs magnetometry was used in conjunction with Hg comagnetometry in the latest upper limit paper [14]. The Cs sensors used by the PSI nEDM collaboration were demonstrated to have sensitivities of 0.75-8 pT/$\sqrt{\mathrm{Hz}}$ and absolute accuracies of 45-90 pT per measurement [20]. However, due to the method by which these sensors are operated they are not magnetically silent (Sec. 2), and can therefore not be used while neutrons are being measured and thus have limited utility.
@@
In this work we demonstrate a statistical sensitivity of 3.5 pT/$\sqrt{\mathrm{Hz}}$ at 1 Hz and a stability of 90 fT over 150 seconds of measurement in a configuration that operates purely optically, using NMOR to both create and probe a polarized atomic state. This allows precision magnetic field measurements to be done during the EDM experiment without perturbing the neutrons [25]. Novel aspects of the design include completely commercial components (Sec. 2) and the ability to run many sensors on a single laser diode (Sec. 4).
@@ 2 Configuration
Figure 1: Shown is a schematic of the vapour cell with the orientation of the magnetic field and optical beams used for this magnetometer. Both the pump and probe beams are linearly polarized with their electric field vectors aligned and perpendicular to the magnetic field direction. The probe beam propagates through the vapour cell along the magnetic field vector to be measured, while the pump beam can be anywhere in the plane formed by the two beams in this figure. The magnetic field information is encoded in the time dependence of the polarization angle $\phi$ of the probe beam after it exits the vapour cell.
@@
The Bell-Bloom style Free Precession Decay (FPD) configuration used in our system also measures the Larmor frequency of Cs to deduce the magnetic field experienced by the atoms; however, it uses nonlinear magneto-optical rotation [26, 29] rather than absorption to both polarize the atoms and interrogate their Larmor frequency. In this configuration, the electric field vector of both the pump and probe beams are perpendicular to the magnetic field, as in figure 1. The linearly polarized light can be described as an equal mixture of left- and right-handed CPL, and so will induce $\mathrm{m}_{\mathrm{F^{\prime}}}=\mathrm{m}_{\mathrm{F}}\pm 1$ transitions. The natural decay of the excited state has a random angular momentum transfer, and so on average the atoms are pushed towards states with $\mathrm{m}_{\mathrm{F}}=\pm$F. Classically this corresponds to the spins of the atoms in the ensemble having a preferred axis of alignment, so this is called an “aligned” state, and the atoms can be said to have been polarized. The magnetic field to be measured now exerts a net torque on these aligned spins, and the atoms precess in the magnetic field at their Larmor frequency. The pump beam is quickly turned off to allow precession to occur. After half of one Larmor period the alignment axis of the atoms is along the laser polarization axis once again, and the pump beam is switched back on, reinforcing the aligned state. By repeating this process, a large population of such polarized atoms can be produced. Once the population is maximized, the pump beam is switched off. A separate, much weaker probe beam is then passed through the atoms, propagating parallel to the magnetic field. The polarized state of the Cs exhibits an axis of birefringence that precesses along with the atoms. This causes the angle of the linearly polarized probe light, $\phi$, to be modulated at twice the Larmor frequency of the atoms because of the twofold symmetry of the aligned state. This modulation is detected via polarimetry after the probe beam exits the vapour.
@@
The FPD configuration eliminates the issue of RF crosstalk and interference by eliminating the use of RF entirely. The issue of phase drift is eliminated by running exclusively in FPD mode, while the issue of the vector light shift is reduced by using linearly polarized light for the probe beam. Additionally, servicing all sensors with a single probe laser means any systematic shifts associated with laser diode intensity or frequency fluctuations should not impact relative measurements between sensors, meaning gradient extraction is not affected. These same changes are planned by the PSI collaboration for their upcoming n2EDM experiment [2] and have been implemented by the panEDM collaboration as well [1, 30, 31, 32]. The sensors used in our work were not made in-house, but rather by an externally contracted commercial manufacturer who assembled the sensors using entirely commercially available components. This project would not have been possible without the involvement of the commercial vendor, Southwest Sciences [33]. Using commercial components allows us to easily expand the system to accommodate more sensors (Sec. 4). The use of commercially purchased components and the decision to permanently epoxy the vapour cells in the sensor heads are the main features that distinguish our system from the panEDM system. Using the BMSR2 at PTB in Berlin the panEDM team was able to produce an extremely stable field using SQUID feedback, and they demonstrated an impressive 600-700 fT average difference between successive measurements [30]. This can most closely be compared to our Allan deviation at minimum integration time, which was 3.9 pT (Sec. 4). They also demonstrated a stability below 50 fT between 70 and 600 s.
@@ 3 Apparatus
All fibre in this system before the sensor is either polarization maintaining (PM) or polarizing (PZ) fibre. This ensures very clean linear polarization in the sensor head. The pump laser is launched into the input fibre of a fibre-coupled acousto-optic modulator (AOM). This is then attached to a $1\times 8$ fibre splitter (also using PM fibre) which splits the pump light equally into 8 pigtailed PM fibres, each of which can provide pump light to a sensor. The probe light is launched directly into another $1\times 8$ PM splitter. The overall intensity of the probe beams are controlled by a linear polarizer and half-wave plate in the free space section of the beam. Individual ports can be additionally attenuated using an adjustment screw on the body of the splitter. The free space beams are $\sim$100 mW, typical launch efficiencies with the current opto-mechanical setup are 20-30%, leading to $\sim$1-2 mW available for use in the sensor heads after accounting for all losses from splitting and attenuation. This is sufficient power to optimize the pump light [25] and more than enough to provide probe power. The pump beams can be globally attenuated by adjusting the high level of the square wave operating the AOM. The probe beams are typically $<5~{}\mu$W in the sensor head after being adjusted using screw-attenuators built into the fibre splitter output ports. $<5~{}\mu$W is enough to reach shot-noise-limited performance in the photodiodes without limiting the coherence time of the Cs via de-pumping. The exact power is optimized to minimize the uncertainty in the frequency fit of the FPD signal.
The magnetically sensitive component of these sensors is an evacuated glass cell with a small amount of Cs held in a reservoir, attached to the cell via a capillary tube. This cell was manufactured by Precision Glass Blowing (PGB) [34]. The cell is a cylinder 30 mm in length and 10 mm in diameter. The cell is treated with an anti-relaxation alkene coating available from PGB prior to being filled, which greatly extends the polarization lifetime of the atoms in the cell by reducing polarization-destroying wall collisions. Prior to being sent to be installed by the contracted manufacturer the important atomic properties of the cell — the coherence time, vapour density etc. — were evaluated at the University of Winnipeg optics lab. It was possible via careful heat treatment and handling to improve these properties. Ultimately just over half of the cells purchased from PGB were determined to be of high enough quality to be installed in sensors. Light is delivered to the Cs cell where it is collimated into a free-space beam to pass through the cell and interact with the Cs. After passing though the Cs cell the probe beam is directed through a polarizing beamsplitter and the resulting two beams are launched into MM fibre to be carried out of the magnetically sensitive environment for detection and analysis. This sensor head was designed and manufactured by Southwest Sciences [33] in consultation with TUCAN, and incorporates design improvements developed in Munich [1, 30, 31, 32], mainly the use of the clamshell design, which will now be described.
@@
The optical components are held in a plastic clamshell with a very similar design to the current panEDM Cs design [30], 3D printed in undyed ABS plastic. The long axis of the cell is transverse to the direction of the incoming fibre. This results in the sensitive axis of the sensor being transverse to the long dimension of the sensor head (Fig. 3). Optical components are cemented to one another at the optical interfaces using RD3-74 epoxy [35]. The epoxy provides the structural connection and maintains alignment between optical components, with the clamshell providing a mounting point for the whole sensor. In the first TUCAN experimental phase these sensors will not be operated in vacuum, so the vacuum compatibility of these materials has not been assessed, however we expect that the components used are vacuum compatible. The optical epoxy ensures efficient light extraction from the sensor head. Linearly polarized light from two laser diodes is delivered to the sensor head via polarizing fibre optics. A gradient-index (GRIN) lens launches light from the probe fibre to a $\sim$1 mm2 free-space beam, which is then directed through a clean-up linear polarizer, followed by the vapour cell via internal reflection through a right-angled prism. This probe launching assembly is mounted to the main ABS body using a coupling which constrains the motion of the probe assembly to linear motion in a plane, which helps with alignment. Linear translation stages are then used to position the assembly to maximize coupling into the exit fibers prior to cementing the prism to the face of the vapour cell as the final step in optical alignment. The beam is then again redirected via a right-angled prism through a Rochon walk-off polarizing beam splitter which splits the beam into orthogonal, linearly polarized components. These components are then launched back into fibre optics: much larger core MM fibres are used on this side so as to collect the most light possible. The two polarization components are then used for polarimetry of the probe light. A GRIN lens also launches linearly polarized pump light into the cell after passing it through a clean-up polarizer, due to the geometry of our setup it can pass straight through the cell at 90 to the probe beam, since the electric field vectors of the two beams are aligned. This choice of beam geometry was largely driven by considering ease of manufacture, launching both beams into the end of the vapour cell would be difficult to do. Any loss in pumping efficiency due to the shorter optical path length can be compensated with pump power.
@@
The Rochon prism is oriented with its fast-axis at a 45 angle relative to the probe beam polarization axis so that it evenly splits the unperturbed probe light into two components of equal power. This means that for small rotations, the degree of rotation of polarized light is directly proportional to the difference in power between the two components. The MM fibres are plugged into a polarimetry board, which amplifies the photo-current difference from the two photodiodes watching the two fibres. The resulting signal is then captured on an oscilloscope, or via a 32 channel DAQ system made by D-tAqc [36]. The D-tAqc system has a bandwidth of 200 kHz with 16 bit resolution. The polarimetry board differencing circuit has a high-frequency cutoff of around 15 kHz, and is AC coupled since we are not interested in static imbalances in the two polarization channels. This relatively low frequency cut-off was due to regulatory requirements imposed on the manufacturer rather than any physics motivation. Figure 4 compares the signal from the board in three states; unpowered, dark, and with the pump beam continuously driving oscillation at the Larmor frequency. Each of these signals is integrated for a full second. The dark noise of the polarimetry board is below the shot-noise limit for the power of the probe beam, meaning the electronics are capable of shot-noise limited detection of the signal.
@@
Figure 6: The amplified, differential photo-current corresponding to the degree of rotation of polarized light as a function of time during the polarization and the free spin precession phases of the Cs gas. Pumping begins at T=0 on this plot, and stops just after 40 ms when the signal reaches its maximum. During this time the pump beam is modulated into a 20% duty cycle square wave at $\sim$7 kHz, approximately the Larmor frequency of Cs in a 1 $\mu$T field. The remainder of the signal is the free spin precession portion which is later fit to an exponentially decaying sine wave. A single instance of this pump-probe cycle is sometimes referred to as one free precession decay cycle, or one “FPD” cycle
Figure 7: The measured magnetic field (blue line) and measured coil voltage (orange line) versus time. The voltage across is being monitored by an 8.5 digit multimeter from Keithley, which is set to an integration time of 4/60 of a second, or 4 power line cycles, matching the integration time of the FPD signal measuring the magnetic field
@@
To operate in 1 $\mu$T we use passive magnetic shielding to reduce both earth’s field and fluctuations in it. The shield is four nested cylinders of $\mu$-metal, a highly magnetically permeable alloy of nickel. The outer cylinder is 40 cm long with a 15 cm radius. Since the sensitive axis of the sensors is transverse to the long axis of the sensor (Fig. 3), the innermost volume contains a 3D printed saddle-wound coil designed to provide a homogeneous transverse field inside the shield. The volume inside this coil is 20 cm long with radius 3.5 cm. The axial shielding factor has been measured [37] to be 1.4×$\times 10^{7}$ with all endcaps installed, and while the transverse shielding factor has not been measured it is expected to be much higher. Since this inner volume is quite small, the uniformity is on the order of nT/cm. This is quite non-uniform, but provides very high AC and DC shielding factors. All sensor tests are done in this shield.
@@ 4 Results
Figure 7 shows a magnetic field measurement with the system. After being split into individual FPDs the signal is filtered with a windowed-sinc FIR filter, with a pass-band from 6-8 kHz, transitional bands of 2 kHz on each side, and $<0.04$ dB of ripple in the pass band. The filtered signal is fit to an exponentially decaying sine wave
@@
where $A$ is the amplitude, $T_{2}$ is the time constant of the decay envelope, $\omega$ is the frequency of the precession, and $\Phi$ is the phase. $\omega$ is directly scaled to be the magnetic field strength via the gyromagnetic ratio of Cs. While this data was taken, the coil voltage was being monitored across a precision resistor using a digital multimeter (DMM) with 8.5 digits of precision. The voltage is scaled to a magnetic field via the ratio of the averages of the two data sets so they can be directly compared without a perfectly accurate coil constant. The Allan deviation of the same data is shown in Figure 8. The voltage monitor from the power supply is significantly less stable than the sensor readings for most integration times. We attribute this to environmental effects on both the DMM itself and the connection to the monitor port. The sensor’s Allan deviation reaches a minimum of 90$\pm$8 fT at an integration time of 150 seconds, so we can conclude that our sensors are at least this stable. It is encouraging that we were able to get within a factor of two of the panEDM stability result [30] despite not having access to the PTB facilities. The minimum integration time Allan deviation corresponds to the average deviation between successive points and can be taken as a measurement of sensitivity. In this case this deviation is 3.9 $\pm$ 0.1 pT. The square root of the power spectrum is a standard measure of sensitivity, for this measurement we find 3.5 pT/$\sqrt{\mathrm{Hz}}$ at 1 Hz.
Figure 9: Magnetic field measurements of two sensors (a and b) together with the difference between them (c) as a function of time. The sensors are operated simultaneously, and when the field is stepped up and down, the change in field is observed in each sensor individually. The sensor measurements are subtracted, as shown in (c), and the difference remains at their combined noise floor. Because of the relatively inhomogeneous field in the magnetic shield, the two sensors are operating at different fields and thus are not able to be perfectly optimized simultaneously given they are driven by the same optical system with a common pump cycle frequency. The higher noise floor is the result of this. The spike in the difference signal around 60 seconds can be associated with the changing magnetic field, and is most likely the result of the field changing partway through an FPD ringdown
@@
This simultaneous operation using a single laser diode with such an easily expandable system is another benefit of the commercial optics were are using. The system as it is set up in figure 5 can operate 8 sensors with no modification whatsoever, provided sufficient magnetically shielded volume, and the TUCAN collaboration has received and successfully operated each of the full compliment of 20 such sensors. The relatively simple addition of more fibre optic splitters will allow the full suite of sensors to be operated once the TUCAN MSR is complete.